# LOLA – An Open-Source Massively Multilingual Large Language Model

### v2

@@ Abstract
Nikit Srivastava, Denis Kuchelev, Tatiana Moteu Ngoli, Kshitij Shetty, Michael Röder, Diego Moussallem, Hamada Zahera, Axel-Cyrille Ngonga Ngomo Paderborn University, Germany nikit.srivastava@uni-paderborn.de
@@ 1 Introduction
Large Language Models (LLMs) have shown tremendous capability across a diverse set of tasks in recent years Radford et al. (2019); Kaddour et al. (2023). This progress has propelled research, with many chat-based LLMs 111 ChatGPT: https://chat.openai.com/; LLAMA: https://llama.meta.com/; Mistral: https://mistral.ai/; Gemini: https://gemini.google.com/; Deepseek: https://www.deepseek.com/. gaining popularity among general users. However, concerns remain, particularly regarding their accessibility for multilingual usage Joshi et al. (2020) and open-source licensing policies Liu et al. (2023). The number of competent LLMs significantly decreases for languages other than English Üstün et al. (2024). This, combined with the curse of multilinguality Conneau et al. (2020), means non-English speakers often have access to inferior systems. Additionally, many new models Jiang et al. (2023); Achiam et al. (2024); Dubey et al. (2024) are pay-to-use, require personal information, or do not fully disclose training details, creating significant hurdles for multilingual research.
To advance multilingual language modeling, we introduce LOLA 222 Github: https://github.com/dice-group/LOLA, HuggingFace: https://hf.co/dice-research/lola_v1 , a massively multilingual model that follows a GPT-style Radford et al. (2019) decoder-only architecture with sparse Mixture-of-Experts (MoE)layers Shazeer et al. (2017). MoE architectures have shown strong performance on learning the underlying structure of the data  Chen et al. (2022) but their application in multilingual LLMs remains underexplored. MoE models can effectively increase model capacity with minimal additional computational cost, offering the possibility of leveraging implicit clusters like language family groups and playing a crucial role in addressing the challenges of multilinguality.
@@
Another important factor is the availability of LLMs as a free resource, accessible for anyone to use, modify, and redistribute without discrimination against any individuals or purposes. Many popular LLMs that claim to be "open source" either withhold their training datasets (e.g., Mistral, Grok [^3] https://github.com/xai-org/grok-1), fail to publish their training code (e.g., Llama, Grok), or do not release their inference code (e.g., Grok-2444https://x.ai/blog/grok-2Spectrum (2024). In some cases, these models are released under licenses that are restrictive, discriminatory, or impose additional conditions Liesenfeld et al. (2023); Liesenfeld and Dingemanse (2024). The artifacts and components used in LOLA were selected based on their suitability for training massively multilingual LLMs while minimizing licensing concerns. All chosen components are obtainable, modifiable, and redistributable in accordance with the terms of their original licenses.
@@ 3 Model Overview
Figure 1: Three-level overview of LOLA architecture. The left-most block provides a high-level overview of the layers within LOLA, including the alternating standard and Mixture-of-Experts (MoE)-based decoder blocks. The middle block gives a detailed view of the MoE-based decoder block structure. The right-most block zooms in on the inner workings of each MoE layer, showing how the top-1 gating mechanism selects from multiple expert FFNs.
Our model is based on a GPT-style Radford et al. (2019) decoder-only Transformer architecture Vaswani et al. (2017). We replace the standard feed-forward layers (FFNs) with Mixture-of-Experts (MoE) layers in every alternate Transformer layer. These MoE layers utilize a top-1 gating mechanism inspired by the Switch Transformer Fedus et al. (2022) due to its simplicity and effectiveness. The architecture consists of 24 decoder layers with a model hidden and embedding dimension of 2048, 16 attention heads, a maximum sequence length of 2048, and each MoE layer includes 16 experts. We use GELU Hendrycks and Gimpel (2017) non-linearities and Adam Kingma and Ba (2015) optimizer for our model. Based on this configuration, our model has 1.3 billion active parameters out of 7.4 billion total parameters. Due to this sparsity, our model has training/inference cost similar to that of a 1.3 billion dense model. [^6] Number of parameters activated in a single forward and backward pass. Figure 1 provides a multi-level overview of the model architecture. The model configuration and training are facilitated using the Megatron-DeepSpeed [^7] https://github.com/microsoft/Megatron-DeepSpeed framework, which is based on Shoeybi et al. (2020); Rajbhandari et al. (2022).
@@ 3.2 Training and Loss Functions
Our model is pre-trained using a causal language modeling task, where the objective is to minimize the cross-entropy loss alongside an auxiliary MoE loss. This auxiliary loss, inspired by works such as Shazeer et al. (2017), Lepikhin et al. (2021), and Fedus et al. (2022), is used to ensure stable training and effective load balancing among experts. Given an input sequence $S=\{s_{1},s_{2},s_{3},\dots,s_{T}\}$ of length $T$, the auxiliary loss $l_{\text{aux}}$ is formulated as
$l_{\text{aux}}=N\cdot\sum_{i=1}^{N}\left(\frac{1}{T}\sum_{t=1}^{T}G_{i}(s_{t})\cdot\frac{1}{T}\sum_{t=1}^{T}M_{i}(s_{t})\right),$ (5)
where $M_{i}(s)$ is a binary mask indicating whether token $s$ is routed to expert $i$, determined by the top-1 gating mechanism.
For the language modeling task, the cross-entropy loss is computed as
$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\sum_{t=1}^{T}\log p(s_{t}|s_{<t}).$ (6)
The final loss function for the model is a combination of the cross-entropy loss ($\mathcal{L}_{\text{CE}}$) and the auxiliary loss ($l_{\text{aux}}$):
@@
where $\alpha$ is the multiplicative coefficient for the auxiliary loss. Throughout this work, we set $\alpha$ to $10^{-2}$ based on the recommendations by Fedus et al. (2022).
@@ 3.3 Training Data and Setup
The model was trained on data sampled from the CulturaX Nguyen et al. (2024a) dataset, which consists of raw text documents in 167 languages, amounting to over 6 trillion tokens from more than 7 billion documents. We tokenized the data using the SentencePiece Kudo and Richardson (2018) tokenizer with a vocabulary size of 100,000.
@@ 4.1 Models
After reviewing the available multilingual LLMs, we selected 17 models with active parameters ranging from 300 million to 7.5 billion. Table 1 provides a list of the selected models along with further details. The selection was based on the following criteria: 1. They are base pretrained models without any fine-tuning; 2. The weights are openly accessible without requiring personal information beyond name and email; 3. Model weights are available via Huggingface [^10] https://huggingface.co/  [^11] Required for the evaluation framework.; 4. The models are compatible with our evaluation hardware setup. [^12] Single NVIDIA A100 with 40GB GPU memory, 100GB of CPU memory, and 16 CPU cores. Given the wide range of active parameters, we decided to group the models based on their sizes. We employ the distortion [^13] Mean sum of squared distances to centers. and silhouette [^14] Mean ratio of intra-cluster and nearest-cluster. scores to determine the optimal number of categories, which was identified as 3 (see Appendix A.1). Subsequently, K-Means clustering was used to classify the models into 3 categories (1-3). Although LOLA falls within Category-1, we compare and analyze it’s performance against each category.
@@ 9 Acknowledgements
This work has been supported by the Ministry of Culture and Science of North Rhine-Westphalia (MKW NRW) within the project SAIL under the grant no NW21-059D; and German Federal Ministry of Education and Research (BMBF) within the EuroStars project E! 114154 PORQUE under the grant no 01QE2056C. The authors gratefully acknowledge the computing time provided to them on the high-performance computer noctua2 Bauer et al. (2024) at the NHR Center PC2. These are funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for the national highperformance computing at universities (www.nhr-verein.de/unsere-partner). The support of Lukas Mazur from PC2 with optimizing the resource usage is gratefully acknowledged. Finally, we extend our sincere gratitude to Dr. Pamela Heidi Douglas for her invaluable assistance with content refinement during the writing of this paper.
@@ A.2 Training Stats
Table 3: Training statistics and model details for LOLA.
@@ A.4 Evaluation Result Tables
We present evaluation results for each model category, as outlined in subsection 4.1. Table 4, Table 5, and Table 6 provide links to the evaluation result tables for Category-1, Category-2, and Category-3, respectively. Additionally, Table 7 contains links to the combined results tables across all categories. Evaluation tables are available at Zenodo.191919 Zero-Shot: https://zenodo.org/records/13750485; One-Shot: https://zenodo.org/records/13750495; Few-Shot: https://zenodo.org/records/13750497.
@@ A.4.1 Missing Results
We acknowledge certain limitations in our experimental setup, particularly where some tables lack results for specific models under certain configurations. [^20] All the files regarding the missing combinations can be found here: https://zenodo.org/records/13763520 We have identified two primary factors contributing to this: 1) Some models (e.g., Glot500m, XLM-R) exhibit limitations in their tokenization logic, which causes them to fail on larger sequences or languages that they cannot tokenize properly; and 2) Certain model architectures (e.g., mT5, mBART, uMT5) are not supported by task implementations such as Belebele. We ensure that the missing results do not negatively impact the comparisons by excluding cases where results are unavailable. Additionally, we take care to prevent these missing results from skewing the visualizations in our favor by accurately representing them as valid gaps.
@@ A.6 Extended MoE Analysis
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption

### v4

@@ Abstract
Nikit Srivastava, Denis Kuchelev, Tatiana Moteu Ngoli, Kshitij Shetty, Michael Röder, Hamada M. Zahera, Diego Moussallem, Axel-Cyrille Ngonga Ngomo Data Science Group, Paderborn University, Germany <firstname.lastname>@upb.de
@@ 1 Introduction
Large Language Models (LLMs) have shown tremendous capability across a diverse set of tasks in recent years Radford et al. (2019); Kaddour et al. (2023). This progress has propelled research, with many chat-based LLMs 111 ChatGPT: chat.openai.com; LLAMA: llama.meta.com; Mistral: mistral.ai; Gemini: gemini.google.com; Deepseek: deepseek.com. gaining popularity among general users. However, concerns remain, particularly regarding their accessibility for multilingual usage Joshi et al. (2020) and open-source licensing policies Liu et al. (2023). The number of competent LLMs significantly decreases for languages other than English Üstün et al. (2024). This, combined with the curse of multilinguality Conneau et al. (2020), means non-English speakers often have access to inferior systems. Additionally, many new models Jiang et al. (2023); Achiam et al. (2024); Dubey et al. (2024) are pay-to-use, require personal information, or do not fully disclose training details, creating significant hurdles for multilingual research.
To advance multilingual language modeling, we introduce LOLA 222 Source Code: github.com/dice-group/LOLA; Model Weights: huggingface.co/dice-research/lola_v1. , a massively multilingual model that follows a GPT-style Radford et al. (2019) decoder-only architecture with sparse Mixture-of-Experts (MoE)layers Shazeer et al. (2017). MoE architectures have shown strong performance on learning the underlying structure of the data  Chen et al. (2022) but their application in multilingual LLMs remains underexplored. MoE models can effectively increase model capacity with minimal additional computational cost, offering the possibility of leveraging implicit clusters like language family groups and playing a crucial role in addressing the challenges of multilinguality.
@@
Another important factor is the availability of LLMs as a free resource, accessible for anyone to use, modify, and redistribute without discrimination against any individuals or purposes. Many popular LLMs that claim to be "open source" either withhold their training datasets (e.g., Mistral, Grok333github.com/xai-org/grok-1), fail to publish their training code (e.g., Llama, Grok), or do not release their inference code (e.g., Grok-2444x.ai/blog/grok-2Spectrum (2024). In some cases, these models are released under licenses that are restrictive, discriminatory, or impose additional conditions Liesenfeld et al. (2023); Liesenfeld and Dingemanse (2024). The artifacts and components used in LOLA were selected based on their suitability for training massively multilingual LLMs while minimizing licensing concerns. All chosen components are obtainable, modifiable, and redistributable in accordance with the terms of their original licenses.
@@ 3 Model Overview
Figure 1: Three-level overview of the LOLA architecture. The left-most block provides a high-level overview of the layers within LOLA, including the alternating standard and Mixture-of-Experts (MoE)-based decoder blocks. The middle block gives a detailed view of the MoE-based decoder block structure. The right-most block zooms in on the inner workings of each MoE layer, showing how the top-1 gating mechanism selects from multiple expert FFNs.
Our model is based on a GPT-style Radford et al. (2019) decoder-only Transformer architecture Vaswani et al. (2017). We replace the standard feed-forward layers (FFNs) with Mixture-of-Experts (MoE) layers in every alternate Transformer layer. These MoE layers utilize a top-1 gating mechanism inspired by the Switch Transformer Fedus et al. (2022) due to its simplicity and effectiveness. The architecture consists of 24 decoder layers with a model hidden and embedding dimension of 2048, 16 attention heads, a maximum sequence length of 2048, and each MoE layer includes 16 experts. We use the GELU Hendrycks and Gimpel (2017) non-linearities and the Adam Kingma and Ba (2015) optimizer for our model. Based on this configuration, our model has 1.3 billion active parameters out of 7.4 billion total parameters. Due to this sparsity, our model has a training/inference cost similar to that of a 1.3 billion dense model. [^6] Number of parameters activated in a single forward and backward pass. Figure 1 provides a multi-level overview of the model architecture. The model configuration and training are facilitated using the Megatron-DeepSpeed777github.com/microsoft/Megatron-DeepSpeed framework, which is based on Shoeybi et al. (2020); Rajbhandari et al. (2022).
@@ 3.2 Training and Loss Functions
Our model is pre-trained using a causal language modeling task Radford and Narasimhan (2018), where the objective is to minimize the cross-entropy loss alongside an auxiliary MoE loss.This auxiliary loss, inspired by works such as Shazeer et al. (2017), Lepikhin et al. (2021), and Fedus et al. (2022), is used to ensure stable training and effective load balancing among the experts. The auxiliary loss incorporates two vectors:
  • $P$ represents the average weight assigned to all tokens for each expert.
  • $f$ denotes the fraction of tokens allocated to each expert.
Given an input sequence $S=\{s_{1},s_{2},s_{3},\dots,s_{T}\}$ of length $T$, and $N$ experts in each MoE layer, for each expert $i=1,2,\dots,N$, these vectors are defined as:
$P_{i}=\frac{1}{T}\cdot\sum_{t=1}^{T}G_{i}(s_{t}),$ (5)
$f_{i}=\frac{1}{T}\cdot\sum_{t=1}^{T}M_{i}(s_{t}),$ (6)
where:
  • $G_{i}(s_{t})$ is the gating weight assigned to expert $i$ for token $s_{t}$,
  • $M_{i}(s_{t})$ is a binary mask indicating whether token $s_{t}$ is routed to expert $i$, determined by the top-1 gating mechanism Shazeer et al. (2017).
The auxiliary loss $l_{\text{aux}}$ is formulated as:
$l_{\text{aux}}=N\cdot\sum_{i=1}^{N}P_{i}\cdot f_{i},$ (7)
which represents the scaled dot product between $P$ and $f$.
For the language modeling task, the cross-entropy loss is computed as:
$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\cdot\sum_{t=1}^{T}\log p(s_{t}\mid s_{<t}),$ (8)
The final loss function for the model combines the cross-entropy loss and the auxiliary loss:
@@
where $\alpha$ is the multiplicative coefficient for the auxiliary loss. Throughout this work, we set $\alpha=10^{-2}$ based on the recommendations by Fedus et al. (2022).
@@ 3.3 Training Data and Setup
The model was trained on data sampled from the CulturaX Nguyen et al. (2024a) dataset, which consists of raw text documents in 167 languages, amounting to over 6 trillion tokens from more than 7 billion documents (see Appendix A.6 for train sample details). We tokenized the data using the SentencePiece Kudo and Richardson (2018) tokenizer with a vocabulary size of 100,000.
@@ 4.1 Models
After reviewing the available multilingual LLMs, we selected 17 models with active parameters ranging from 300 million to 7.5 billion. Table 1 provides a list of the selected models along with further details. The selection was based on the following criteria: 1. They are base pretrained models without any fine-tuning; 2. The weights are openly accessible without requiring personal information beyond name and email; 3. Model weights are available via Huggingface101010huggingface.co  [^11] Required for the evaluation framework.; 4. The models are compatible with our evaluation hardware setup. [^12] Single NVIDIA A100 with 40GB GPU memory, 100GB of CPU memory, and 16 CPU cores. Given the wide range of active parameters, we decided to group the models based on their sizes. We employ the distortion [^13] Mean sum of squared distances to centers. and silhouette [^14] Mean ratio of intra-cluster and nearest-cluster. scores to determine the optimal number of categories, which was identified as 3 (see Appendix A.1). Subsequently, K-Means clustering was used to classify the models into 3 categories (1-3). Although LOLA falls within Category-1, we compare and analyze its performance against each category.
@@ 9 Acknowledgements
This work has been supported by the Ministry of Culture and Science of North Rhine-Westphalia (MKW NRW) within the project SAIL under the grant no NW21-059D; and German Federal Ministry of Education and Research (BMBF) within the EuroStars project E! 114154 PORQUE under the grant no 01QE2056C. The authors gratefully acknowledge the computing time provided to them on the high-performance computer noctua2 Bauer et al. (2024) at the NHR Center PC2. These are funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for the national high-performance computing at universities (www.nhr-verein.de/unsere-partner). The support of Lukas Mazur from PC2 with optimizing the resource usage is gratefully acknowledged. Finally, we extend our sincere gratitude to Dr. Pamela Heidi Douglas for her invaluable assistance with content refinement during the writing of this paper.
@@ A.2 Training Stats
Table 3: Training statistics and model details for LOLA.
@@ A.4 Evaluation Result Tables
We present evaluation results for each model category, as outlined in subsection 4.1. Table 4, Table 5, and Table 6 provide links to the evaluation result tables for Category-1, Category-2, and Category-3, respectively. Additionally, Table 7 contains links to the combined results tables across all categories. Evaluation tables are available at Zenodo.191919 Zero-Shot: zenodo.org/records/13750485; One-Shot: zenodo.org/records/13750495; Few-Shot: zenodo.org/records/13750497.
@@ A.4.1 Missing Results
We acknowledge certain limitations in our experimental setup, particularly where some tables lack results for specific models under certain configurations. [^20] All the files regarding the missing combinations can be found here: zenodo.org/records/13763520 We have identified two primary factors contributing to this: 1) Some models (e.g., Glot500m, XLM-R) exhibit limitations in their tokenization logic, which causes them to fail on larger sequences or languages that they cannot tokenize properly; and 2) Certain model architectures (e.g., mT5, mBART, uMT5) are not supported by task implementations such as Belebele. We ensure that the missing results do not negatively impact the comparisons by excluding cases where results are unavailable. Additionally, we take care to prevent these missing results from skewing the visualizations in our favor by accurately representing them as valid gaps.
@@ A.6 Extended MoE Analysis
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
Code Language Doc. Count
af Afrikaans 33 0603306033\,06033 060
als Swiss German $6936$
am Amharic $9733$
an Aragonese $2746$
ar Arabic 2 961 11829611182\,961\,1182 961 118
arz Egyptian Arabic 71 6257162571\,62571 625
as Assamese 52 6275262752\,62752 627
ast Asturian $9002$
av Avaric $438$
az Azerbaijani 203 380203380203\,380203 380
azb South Azerbaijani 29 8332983329\,83329 833
ba Bashkir 71 9577195771\,95771 957
bar Bavarian $3$
bcl Central Bikol $1$
be Belarusian 65 7396573965\,73965 739
bg Bulgarian 965 272965272965\,272965 272
bh Bihari languages $265$
bn Bangla 497 463497463497\,463497 463
bo Tibetan 54 1855418554\,18554 185
bpy Bishnupriya $5087$
br Breton 43 7654376543\,76543 765
bs Bosnian $1237$
bxr Russia Buriat $100$
ca Catalan 621 271621271621\,271621 271
cbk Chavacano $2$
ce Chechen 17 3221732217\,32217 322
ceb Cebuano 10 5551055510\,55510 555
ckb Central Kurdish $6881$
cs Czech 2 614 02226140222\,614\,0222 614 022
cv Chuvash 22 5702257022\,57022 570
cy Welsh 21 9982199821\,99821 998
da Danish 1 017 19210171921\,017\,1921 017 192
de German 67 202 7976720279767\,202\,79767 202 797
dsb Lower Sorbian $59$
dv Divehi 66 7026670266\,70266 702
el Greek 2 057 20920572092\,057\,2092 057 209
eml Emiliano-Romagnol $91$
en English 64 821 3136482131364\,821\,31364 821 313
eo Esperanto 18 4031840318\,40318 403
es Spanish 18 037 5051803750518\,037\,50518 037 505
et Estonian 320 190320190320\,190320 190
eu Basque 63 9526395263\,95263 952
fa Persian 2 381 24523812452\,381\,2452 381 245
fi Finnish 1 218 70612187061\,218\,7061 218 706
fr French 14 550 1731455017314\,550\,17314 550 173
frr Northern Frisian $11$
fy Western Frisian $8930$
ga Irish 12 1701217012\,17012 170
gd Scottish Gaelic $8408$
gl Galician 71 4387143871\,43871 438
gn Guarani $103$
gom Goan Konkani $721$
gu Gujarati 46 5154651546\,51546 515
he Hebrew 186 159186159186\,159186 159
hi Hindi 786 614786614786\,614786 614
hr Croatian 18 4271842718\,42718 427
hsb Upper Sorbian $4244$
ht Haitian Creole $12$
hu Hungarian 1 765 28617652861\,765\,2861 765 286
hy Armenian 118 579118579118\,579118 579
ia Interlingua $613$
id Indonesian 930 054930054930\,054930 054
ie Interlingue $4$
ilo Iloko $2328$
io Ido $1144$
is Icelandic 94 9429494294\,94294 942
it Italian 8 452 39684523968\,452\,3968 452 396
ja Japanese 4 447 53944475394\,447\,5394 447 539
jbo Lojban $1349$
jv Javanese $2058$
ka Georgian 124 812124812124\,812124 812
kk Kazakh 109 359109359109\,359109 359
km Khmer 40 5274052740\,52740 527
kn Kannada 54 0855408554\,08554 085
ko Korean 822 292822292822\,292822 292
krc Karachay-Balkar $1745$
ku Kurdish 11 8121181211\,81211 812
kv Komi $1396$
kw Cornish $94$
ky Kyrgyz 22 8362283622\,83622 836
la Latin 48 9684896848\,96848 968
lb Luxembourgish $6635$
lez Lezghian $1806$
li Limburgish $206$
lmo Lombard $3530$
lo Lao $8713$
lrc Northern Luri $43$
lt Lithuanian 533 591533591533\,591533 591
lv Latvian 285 463285463285\,463285 463
mai Maithili $93$
mg Malagasy $4636$
mhr Eastern Mari $7883$
min Minangkabau $1429$
mk Macedonian 110 512110512110\,512110 512
ml Malayalam 107 722107722107\,722107 722
mn Mongolian 77 1537715377\,15377 153
mr Marathi 90 6639066390\,66390 663
mrj Western Mari $1056$
ms Malay $9526$
mt Maltese $6052$
mwl Mirandese $9$
my Burmese 34 6233462334\,62334 623
myv Erzya $4$
mzn Mazanderani $1914$
nah Nahuatl languages $131$
nap Neapolitan $31$
nds Low German 15 1391513915\,13915 139
ne Nepali 124 961124961124\,961124 961
new Newari $4344$
nl Dutch 4 695 70646957064\,695\,7064 695 706
nn Norwegian Nynorsk $5043$
no Norwegian 756 292756292756\,292756 292
oc Occitan 10 5561055610\,55610 556
or Odia $6138$
os Ossetic $8596$
pa Punjabi 25 8792587925\,87925 879
pam Pampanga $4$
pl Polish 5 686 68856866885\,686\,6885 686 688
pms Piedmontese $7566$
pnb Western Panjabi 15 6251562515\,62515 625
ps Pashto 15 0761507615\,07615 076
pt Portuguese 7 611 58676115867\,611\,5867 611 586
qu Quechua $1202$
rm Romansh $30$
ro Romanian 1 613 01616130161\,613\,0161 613 016
ru Russian 31 972 4363197243631\,972\,43631 972 436
rue Rusyn $1$
sa Sanskrit 16 2901629016\,29016 290
sah Sakha 22 1412214122\,14122 141
scn Sicilian $21$
sd Sindhi $4366$
sh Serbian (Latin) 45 6194561945\,61945 619
si Sinhala 30 1463014630\,14630 146
sk Slovak 743 300743300743\,300743 300
sl Slovenian 293 415293415293\,415293 415
so Somali $39$
sq Albanian 208 223208223208\,223208 223
sr Serbian 162 126162126162\,126162 126
su Sundanese $1554$
sv Swedish 1 988 36719883671\,988\,3671 988 367
sw Swahili 66 5066650666\,50666 506
ta Tamil 189 138189138189\,138189 138
te Telugu 72 9147291472\,91472 914
tg Tajik 19 3531935319\,35319 353
th Thai 838 422838422838\,422838 422
tk Turkmen 14 3931439314\,39314 393
tl Filipino 13 9381393813\,93813 938
tr Turkish 3 768 29837682983\,768\,2983 768 298
tt Tatar $8724$
tyv Tuvinian $23$
ug Uyghur 47 0354703547\,03547 035
uk Ukrainian 1 789 62117896211\,789\,6211 789 621
ur Urdu 110 291110291110\,291110 291
uz Uzbek 87 2198721987\,21987 219
vec Venetian $113$
vi Vietnamese 4 096 44740964474\,096\,4474 096 447
vls West Flemish $1$
vo Volapük $6621$
wa Walloon $1383$
war Waray 23 6872368723\,68723 687
wuu Wu Chinese $222$
xal Kalmyk $51$
xmf Mingrelian $9706$
yi Yiddish $5646$
yo Yoruba $192$
yue Yue Chinese $3$
zh Chinese 8 744 98487449848\,744\,9848 744 984
Total Doc. Count 275 653 546275653546275\,653\,546275 653 546
Table 8: List of languages included in the CulturaX dataset, along with the corresponding number of documents per language in the training sample used for LOLA. The language codes utilized are derived from the CulturaX dataset, which adheres to a combination of ISO 639-1 and ISO 639-3 standards. An exception is the use of als, which is considered obsolete; the ISO 639-3 standard designates gsw as its replacement.